OpenStudy (anonymous):

what is the relative minimum and maximum of f(x)= x + 1/x

OpenStudy (anonymous):

You need derive the function with respect to x after igual a cero y will obtain min{f(x)}=-1 max{f(x)}=1

OpenStudy (anonymous):

ok, well when i derived it i got f(x)= 2/x^3. is that correct? and how do you tell what the relative min and max are from there?

OpenStudy (anonymous):

\[df(x)/dx=(dx/dx=1)+(d(1/x)/dx=-1/x^2)\] so that \[df(x)/dx=1-1/x^2\] after \[df(x)/dx=1-1/x^2=0\] x=1 x=-1

OpenStudy (anonymous):

Hope that helps